OpenStudy (anonymous):

Solve the proportion 35/28=x/4 . What is the value of x

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

Sorry, so x = 5.

OpenStudy (hoblos):

multiply both sides by 4 => x=140/28=5

OpenStudy (anonymous):

Thank you :) To solve this, do I take the reciprocal of 35/28 and multiply by 4??

OpenStudy (hoblos):

no just multiply by 4

OpenStudy (anonymous):

that is correct, but just in case lay it out like this. cross multily 28x=140 then divide to get x by it's self x=5

OpenStudy (anonymous):

You can do =) Or divide 35/28 then multiply that by. I just did it mentally - 28/4 = 7, 35/7 = 5.

OpenStudy (anonymous):

oh okay :) you guys are awesome thank u